### v1

# Emergent Hierarchy in Localized States of Organic Quantum Chains

Abstract
Organic Quantum Chains (OQCs) represent a newly synthesized class of carbon-based nanostructures whose quasi-one-dimensional nature gives rise to unconventional electronic and transport phenomena. Here we investigate the electronic and transport properties of recently synthesized OQCs [Nature Communications, 12, 5895 (2021)]. Structural stability was first assessed through molecular dynamics relaxation combined with density functional theory (DFT). The optimized coordinates are then used in a tight-binding model with exponentially decaying hopping parameterization, which reproduces the DFT results with high accuracy. Our calculations reveal a robust and nearly constant energy gap across several OQC configurations, in agreement with experimental data. We also identify emergent hierarchical states, characterized by distinct localization behaviors within sets of localized bands. Finally, we analyze different transport responses in scenarios involving the one-dimensional OQC coupled to carbon corrals, as observed in the experimental data, highlighting their potential as promising systems for application in carbon nanodevices.
keywords:
Organic quantum chains , Carbon nanostructures , Density Functional Theory , Tight-Binding model , Quantum transport , Localized states
[…]
[uff]organization=Instituto de Física, Universidade Federal Fluminense, addressline=Av. Litorânea s/n, city=Niterói, postcode=24210-340, state=RJ, country=Brazil
[…]
[unicamp]organization=Gleb Wataghin Institute of Physics, Department of Applied Physics, State University of Campinas, addressline=Rua Sérgio Buarque de Holanda, 777, city=Campinas, postcode=13083-859, state=SP, country=Brazil
[…]
[unb]organization=Department of Electrical Engineering, College of Technology, University of Brasília, city=Brasília, postcode=70910-900, state=DF, country=Brazil
1 Introduction
Organic Quantum Chains (OQC) are a class of carbon-based systems that exhibit peculiar electronic and structural properties due to their quasi-one-dimensional architecture [1, 2, 3, 4]. These systems are composed of conjugated organic molecules arranged in a chain-like fashion, enabling efficient charge transport and tunable electronic behavior. The properties of OQCs make them promising candidates for applications in molecular electronics, optoelectronics, and quantum-spin computing [5, 6, 7, 4, 8, 9]. Such low-dimensional systems have been synthesized employing self-assembly within chemical routes and direct manipulation with scanning microscope tips [10]. Moreover, when constructed exclusively from carbon-based molecular building blocks (e.g., polycyclic aromatic hydrocarbons, graphene nanoribbons, or conjugated polymers), these systems exhibit emergent electronic localization, topological, and magnetic properties absent in bulk materials [11, 12, 13, 14]. Very recently, nanoscaled chains were synthesized during a growth process of organic molecules on a gold substrate [1].
Figure 1: (a) Experimental STM image highlighting distinct motifs: ring-like structures (red arrow) and chain segments (green arrow) [Adapted [1]]. (b) Molecular precursor used in the synthesis. (c) Schematic representation of the pairing process that connects two precursors. (d) Resulting N1-OQC, representing the smallest unit cell chain segment. (e-m) Optimized geometries of Ni-OQCs for $i=2$ to $10$, illustrating the structural evolution and increasing chain complexity along the periodic direction ($L_{\infty}$).
Stable macrocycle quantum corrals composed of twelve molecular units, marked with a red arrow in Fig. 1(b), are also generated, providing new designed quantum states. Such structures have a precursor unit composed of four hexagons and one pentagon, as depicted in Fig. 1(b). In the sequence, Fig. 1(c) shows the building pair process in the construction of an OQC, exemplified in Fig. 1(d) as an N1-OQC. In this work, we employ various theoretical frameworks for modeling OQCs, with a focus on their electronic structure, quantum transport properties, and potential applications. By combining density functional theory (DFT) with Green’s function (GF) formalism, we provide a detailed understanding of the fundamental properties of OQC and their implications for next-generation molecular devices.
Low-dimensional systems are known to exhibit localized states. A distinct class of localized states, known as compact localized states (CLS), is found in certain low-dimensional periodic systems, such as specific atomic lattices or photonic crystals [11, 15, 16]. Unlike typical wavefunction density probabilities that either fully extend or decay exponentially in space, a CLS has a non-zero amplitude only within a strictly finite and small region, abruptly becoming zero everywhere else. Rather than being caused by randomness or disorder, this extreme form of localization is an intrinsic property of the system’s structure, arising from destructive quantum interference dictated by the specific lattice geometry. The presence of these states is a signature of systems hosting flat bands in their energy spectrum, where the group velocity of particles is zero, effectively trapping them in space and preventing their propagation. Kagomé, Lieb, and fractal lattices, or graphene-based chain lattices [17, 12, 18], are examples of specific geometries that lead to flat bands. The zero-low band dispersion leads to an elevated effective mass of charge holders, obstructing transport and promoting the appearance of topological states of matter, mainly in confined carbon-based structures [19]. Understanding localized states is crucial in designing new devices that harness the robustness and quantum interference of these unique states.
Typically, the coupling of finite molecules to continuum systems results in Fano antiresonance phenomena [20, 21, 22], characterized by a dip in transmission responses due to a destructive interference. On the other hand, in the cases where the molecule states are completely decoupled from the chain system, there is a formation of a perfectly localized mode inside of the continuum state, named a bound quantum state in the continuum (BIC), with no dip on the conductance value related to the decoupled system [23, 24]. BICs are also discussed in wave problems such as acoustics, microwaves and nanophotonics [25, 26, 27]. In this work, we have investigated the emergence of localized states of different natures, creating a unique set of properties that can be directly observed in electronic responses.
2 Theory & Computational details
2.1 Systems Modeling
[…]
According to the experimental report, the precursor molecule tends to exhibit natural curvature, a feature exploited in the context of electronic confinement in corral-like structures [1]. Based on this behavior, a molecular intercalation strategy was adopted to enable the formation of periodic one-dimensional systems with enhanced geometric stability. In this approach, Ni-OQC is defined as a system whose unit cell contains a pair of $i$ covalently connected molecules arranged in sequence. This configuration promotes mutual compensation of local curvature between adjacent molecular chains of the same length, contributing to the overall structural stability. In the present study, values of $i$ ranging from 1 to 10 were investigated. The number of carbon sites $(N)$ in the unit cell is given by $N=40i$, where $i$ is the OQC index. The corresponding unit cells contained between 60 and 600 atoms (carbon and hydrogen), depending on the number of molecules.
The initial geometries were constructed in a linear configuration and did not account for the natural curvature of the OQCs. Given the high computational cost of structural optimization using DFT, an intermediate relaxation stage was implemented using classical molecular dynamics (CMD), carried out with the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [28]. The structural preparation protocol was divided into five stages to generate classically optimized geometries, which were then used as input for subsequent DFT calculations.
The first stage involved the simultaneous minimization of potential energy and internal forces, excluding thermal effects. The total energy was minimized using the conjugate gradient algorithm. The cell volume was also relaxed under zero pressure in the $z$ direction. To escape from local energy minima, a simulation in the NPT ensemble was performed for 50ps at a temperature of 300K and zero pressure along the longitudinal direction. This step induced realistic structural curvatures along the molecular chains. The system was then linearly cooled from 300K to 1K over 25ps in the same ensemble. An additional simulation of 25ps was carried out at a constant temperature of 1K to eliminate residual thermal fluctuations. Finally, a new energy minimization stage was conducted using the same parameters as in the initial stage to ensure geometric stability before DFT optimization.
Throughout the simulation process, Newton’s equations of motion were integrated using the Velocity-Verlet algorithm [29] with a time step of 0.1fs. Interactions between carbon and hydrogen atoms were described using the Adaptive Intermolecular Reactive Empirical Bond Order (AIREBO) potential [30]. Temperature control was implemented through the Nosé-Hoover thermostat [31, 32].
To simulate the interaction between OQCs and a metallic substrate, as occurs in the experimental synthesis, the same systems previously relaxed in the gas phase were repositioned at a distance of 3 Å from the edge of a modified simulation box. This new configuration included non-periodic boundaries and a static substrate modeled through a 12-6 Lennard-Jones (LJ) potential [33], given by
$E_{\text{C-Au}}=4\varepsilon_{\text{C-Au}}\left[\left(\frac{\sigma_{\text{C-Au}}}{r}\right)^{12}-\left(\frac{\sigma_{\text{C-Au}}}{r}\right)^{6}\right],$ (1)
where $r$ is the distance between each atom and the substrate, $\varepsilon_{\text{C-Au}}$ is the potential well depth, and $\sigma_{\text{C-Au}}$ is the equilibrium distance. The interaction is truncated at a cutoff radius $r_{C}$, such that $E_{\text{C-Au}}=0$ for $r>r_{C}$. For the interaction between carbon and gold, we used $\varepsilon_{\text{C-Au}}=\text{31.81}$ meV and $\sigma_{\text{C-Au}}=\text{2.99}$ Å [34]. The interaction between hydrogen and gold was assumed to be negligible. After the classical simulations, the optimized OQC geometries were subjected to a new structural refinement via DFT, as described in the next section.
2.2 Density Functional Theory
The geometric optimization and electronic structure of the OQCs were determined using first-principles calculations based on DFT, implemented in the SIESTA code. This code was selected for its efficiency in handling large-scale systems. The structural optimization was conducted within the generalized gradient approximation (GGA), using the Perdew-Burke-Ernzerhof (PBE) functional, widely recognized for its balance between computational cost and accuracy in describing atomic structures.
To mitigate the well-known underestimation of the band gap associated with GGA functionals [35], we performed additional calculations with the hybrid functional HSE06 [36, 37], implemented through the HONPAS package [38, 39]. Brillouin zone integration was performed with a Monkhorst-Pack grid [40] $1\times 1\times 10$ k along the periodic direction of the chains ($x$). The electron-ion interactions were described by norm-conserved Troullier-Martins pseudopotentials [41, 42], in the factorized Kleinman-Bylander form [43].
We used a 400 Ry grid cutoff for the kinetic energy and a single-$\zeta$ (SZ) basis. Structural optimizations involved complete relaxation of atomic positions and lattice vectors, with convergence criteria of strength of 0.01 eV/Å and energy of $10^{-4}$ eV. To avoid spurious interactions between periodic images, periodic boundary conditions were applied exclusively in the axial direction of the chains, while vacuum regions larger than 35 Åand 50 Åwere established in the transverse ($x$) and normal ($y$) directions, respectively.
2.3 Tight-binding
The interplay of conjugation, symmetry, and quantum confinement effects governs the electronic properties of OQC systems. The quasi-one-dimensional nature of OQC leads to the formation of single-quantum-channel transport along the chain. The OQCs can be described using the tight-binding Hamiltonian,
$H=\sum_{i}\epsilon_{i}c_{i}^{\dagger}c_{i}+\sum_{i,j}t_{ij}\left(c_{i}^{\dagger}c_{j}+\text{H.c.}\right),$ (2)
where $\epsilon_{i}$ represents the on-site energy, $t_{ij}$ is the hopping integral between sites $i$ and $j$, and $c_{i}^{\dagger}$ ($c_{i}$) is the creation (annihilation) operator for electrons at site $i$.
Quantum transport properties of OQCs are investigated using the Green function formalism. The Green function $G(E)$ is defined as
$G(E)=\left[EI-H-\Sigma_{L}(E)-\Sigma_{R}(E)\right]^{-1},$ (3)
where $E$ is the energy and $\Sigma_{L}(E)$ and $\Sigma_{R}(E)$ are the self-energies of the left and right electrodes, respectively. The total density of states (DOS) is obtained by $\text{DOS}=-(1/\pi)\text{Im}\{Tr[G(E)]\}$. The transmission function $T(E)$ is given by [44],
$\mathcal{T}(E)=\text{Tr}\left[\Gamma_{L}(E)G(E)\Gamma_{R}(E)G^{\dagger}(E)\right],$ (4)
where $\Gamma_{L}(E)$ and $\Gamma_{R}(E)$ are the broadening matrices of the left and right electrodes, respectively. To pursue the spatial distribution of electronic transmission, we employ a local transmission coefficient $T^{\alpha}_{ij}(E)$ which computes the transmission between $j$ and $i$ sites, and leads $\alpha=L$ or $R$ [45], namely
$T^{\alpha}_{ij}(E)=2\text{Im}\{[G(E)\Gamma_{\alpha}G^{\dagger}(E)]_{ji}H_{ij}\},$ (5)
Local transmission provides the spatial information of the preferential paths for electrons in a given energy state. A high density in the local transmission implies in a higher probability of the state percolation across the atomic bonds.
The tight-binding parameters used in the whole calculations are obtained by fitting the electronic band structures with our DFT results. We employ the exponential decaying hopping parametrization, given by
$t^{\prime}_{ij}=t_{1}e^{-\beta\big(\frac{r_{ij}}{d_{\text{min}}}-1\big)},$ (6)
with $r_{ij}$ being the distance between $i,j$ lattice sites, $t_{1}$ the hopping related to the first nearest-neighbor distance $d_{\text{min}}=\text{1.37}$ Å, and $\beta$ a fitting parameter that controls the range of the hopping energies. As the ratio $r_{ij}/d_{\text{min}}$ is always larger than one beyond the first nearest-neighbors, small values of $\beta$ increase the number of neighbors with non-negligible hopping contributions in the description. The nearest-neighbor distance between atoms inside the unit cell (black bounds in Fig. 1) is within the range 1.37–1.50 Å, obtained from the relaxed structure. This parametrization scheme has been successfully employed for carbon-based mixed geometries in graphene allotropes, such as multidimensional biphenylene systems [46].
For the chains, the Green function formalism is employed, utilizing the decimation method to achieve numerical periodicity in momentum space. However, for finite systems, such as carbon molecule rings, the local density of states (LDOS) is calculated as a function of energy by
$LDOS(x,y,E)=\sum_{n}|\psi_{n}(x,y)|^{2}\delta(E-E_{n}),$ (7)
where $x$ and $y$ are site coordinates and $\psi_{n}(x,y)$ is the corresponding electronic wave function of the $n^{th}$ state. To compute the delocalized electronic contribution across the system, given by $|\psi_{n}(x,y)|^{2}$, we transform the delta functions into Lorentzian functions, with $\Gamma$ fitted with the experimental results [1].
3 Electronic & Transport Properties
3.1 Organic Quantum Chains (OQCs)
The systems analyzed in this study exhibit distinct geometrical characteristics arising from the intrinsic flexibility of their carbon-hydrogen chains and the resulting curvature. The STM image shown in Fig. 1(a) reveals a variety of configurations, including curved and closed structures, as well as linear segments that connect repeating units aligned along the surface. The internal cavities formed by the spatial arrangement of these units vary in size, leading to a porous and directionally anisotropic morphology.
Figures 1(d-m) present the morphological evolution of the OQC structures as a function of the size of building blocks $N_{i}$, culminating in extended architectures. In particular, the configuration denoted as $N_{10}$ [see Fig. 1(m)] exhibits lateral continuity along the $z$ direction, with a characteristic length $L_{\infty}$ that indicates a periodic vertical arrangement. Despite local variations in cavity size and curvature, the system maintains a well-defined periodicity along the axis of growth.
Figure 2: Electronic band structure evolution of the OQC chains for N(1-10). Red curves are the TB results, black and blue the DFT calculations adopting HSE06 and PBE, respectively. The central energy gap is $\Delta E=\text{2.0}$ eV.
Figure 3: (a) Conductance $G(2e^{2}/h)$] and DOS (a.u.) hierarchy of several OQCs for an arbitrary energy range. A pair of plateaus are separated by a gray line. The gray dashed line separates the left and right leads. Plateaus length in eV for (b) Left and (c) Right panels corresponding to the plateaus positioned at left and right of the gray axis, labeled accordingly panel (a). An exponential decaying fit is displayed at the right of panels (b) and (c). Solid black, dashed gray and black, and solid gray lines, respectively, for 1-4 labeled plateaus panel (a).
Building on this structural understanding and based on the energetic analysis detailed in the A, we now examine how the electronic properties evolve along with the $N_{i}$ cases here explored. Fig. 2 shows the evolution of the TB band structures of the system, with an increasing number of precursors inserted in the unit cell of the OQCs. In the first panel ($N_{1}$-OQC), the results are compared between the TB method in red, and DFT, with PBE and HSE06 functionals; blue and black curves, respectively. The best TB parametrization adopted for all $N_{i}$-OQC structures in the present work are the following; $\beta=2.8/d_{\text{min}}=2.04$, $\epsilon=\text{-0.22}$ eV, and $t_{1}=\text{-3.4}$ eV. The exponentially decaying hopping was limited by a cutoff radius of R<4.2R<{4.2} Å.
As expected, the increasing number of precursors in the unit cell (from $N_{1}$ to $N_{10}$) creates new localized states which may be associated with quasi-finite systems. The gap centered at $E=0$ (Fermi energy) is protected against increasing $N_{i}$, remaining the same ($\Delta E\approx\text{2.0}$ eV). This feature is found in both the DFT and TB results. A zigzag-like pattern inside the band structure is observed as the size of the OQC unit cell increases. This effect resembles the folding effect, i.e., when the system is described by a larger unit cell than necessary; however, each of those bands has separate contributions to the electronic properties. The van Hove singularities in the OQCs density of states follow the band structure hierarchy. We further analyze those patterns by choosing a specific energy range where the band separations lead to the formation of mini-bands, as shown in Fig. 2.
The ballistic conductance plateaus shown in Fig. 3(a) are subdivided into compatible sets, compared to the number of precursors inserted into the unit cell. The gray dashed line separates the two groups of plateaus, left and the right, studied with similar subdivision pattern. The hierarchical structure observed in the conductance sequences is also found for other energy ranges, with similar phenomena. The pseudo-gaps around such finite-like states evolve with new states following the hierarchy of increasing number of precursors in the OQC unit cell.
In order to highlight the hierarchical sequence of conductance plateaus, we show in Fig. 3(b) and (c) an analysis between the plateau length (in eV) according to the labels in Fig. 3(a). The both panels (b) and (c) shows a similar trend between the evolution on the left and right features for several OQC sizes following the colored legend for $N_{1}-N_{9}$. In the right side of panels (b) and (c), respectively, we show the exponential behavior of the plateau length (in eV), considering plateaus labeled from 1 to 4, in solid black, dashed gray, dashed black, and solid gray, respectively, for OQCs $N_{1}-N_{9}$. The overall trend shows a slowing exponential decreasing plateau length as the OQC size changes. In particular, the decreasing factor is attenuated when considering the most internal plateaus (labels >2>2). This behavior is evidenced by the exponential decay of the widths associated with these plateaus, which for increasingly larger building blocks indicates a tendency towards electronic localization.
Figure 4: (a) Localization map for the $N_{4}$ OQC chain, plotting charge density ($|\psi|^{2}$) versus site position for all energy states, with detached zoom for energies $E_{a}$, $E_{b}$, $E_{c}$, and $E_{d}$. The lower panel highlights distinct localization behaviors inside the unit cell. (b) Electronic conductance $G$ (in units of $2e^{2}/h$) and density of states (DOS) compared with the electronic bands.
An overall view of the electronic amplitude, $|\psi_{n}(x,y)|^{2}$, gives a fingerprint of the electronic behavior for the energy states and may serve to distinguish between compact localized and spread (or bulk-like) states. The localization pattern is experimentally feasible, as performed for nanographene-based chains [4]. The heatmap in Fig. 4(a) sheds light on the different features of the states by comparing the corresponding renormalized local density (LDOS); purple (yellow) colored range indicates the less (most) probable charge accumulation at the atomic sites displayed in the x-coordinates (Å). For the $N_{4}$ OQC, four energy values were used as examples $E_{a}=\text{-2.18 }$ eV, $E_{b}=\text{-2.20 }$ eV, $E_{c}=\text{6.51 }$ eV, and $E_{d}=\text{6.60 }$ eV] to show the emergence of different kinds of localized states, being conductive like $E_{b}$, or constricted (without transport responses), as $E_{a,c,d}$. This feature was also verified for different OQCs sizes. We see that state $E_{b}$ has a long-spread density probability that crosses the organic chain unit cell. The energy band in this case has a very small dispersion, giving rise to the conductance channel shown in Fig. 4(b). The energies $E_{a}$ and $E_{c,d}$ correspond to localized states at the edges of the unit cell. They do not allow conductance channels. However, $E_{a}$ has localized states spread over a larger range than in the $E_{c,d}$ cases. This can be seen in the lower panel of Fig. 4(a), where the localization tail goes to zero, being localized in fewer sites, with compact behavior, for $E_{c,d}$. Both $E_{a}$ and $E_{c}$ have edge-like localized states, with no percolation between them, and $E_{d}$ has a charge concentration in the center of the unit cell.
The spreading states exhibit an overall uniform charge distribution over the space, as can be seen for several energies in the complete spectrum of Fig. 4(a). These findings show that OQC systems, with their rich electronic behavior, hold significant promise for carbon-based nanodevice applications.
3.2 Coupled Corrals to OQCs
Apart from the isolated organic corral and chains exhibited in the experimental output shown in Fig. 1(a), coupled quantum corrals and chains are also found in different configurations. Here, we investigate the transport and electronic responses of such systems. Most of the structures observed have $N_{1}$ chains coupled to a quantum ring by a single or a small set of precursors. Here we define the system ring + precursor as “molecule”. We start then by comparing the total density of states (DOS) for a single quantum corral and the molecule system, with the reported [1] experimental results of $dI/dV$ curves, as shown in Figs. 5(a,b). The used tight-binding parametrization is the same as discussed before for the chains. We notice that the calculated DOS (gray curves) agree well with the experimental data. Moreover, the quantum corral has an experimental gap $\Delta E=2.35$ eV, while the theoretical model (TB and HSE06 functional) predicts $\Delta E\approx 2.0$ eV. The experimental DOS results exhibit in-gap states (dashed curve in part (a)) due to the gold substrate, absent in the theoretical description.
The energy states of these corral systems are sensitive to atomic coupled structures, as depicted in Fig. 5(b). The energy spectra of the quantum ring, with and without the precursor, are nearly identical, except for the emergence of electronic states within the pseudo-gaps of the corral system, as highlighted by the blue-shaded regions. In contrast, the coupling of different ions at positions P1, P2, and P3 (magenta, green, and yellow disks in Fig. 5(b)) introduces new atomic states not only within the main energy gap (shaded yellow region) but also in other pseudo-gaps, as indicated by the arrows. This sensitivity demonstrates the potential of coupled quantum corrals as nanoscale sensors in advanced device applications.
Figure 5: (a) Left: Experimental dI/dV spectrum of the quantum ring, showing a gap of $\Delta E=2.35$ eV (yellow). Right: Corresponding calculated density of states (DOS). [Adapted from Ref. [1]]. (b) Comparison of electronic properties for the molecule with and without the precursor. The precursor introduces states (blue) inside the original ring’s pseudo-gaps.
Figure 6: (a) Schematics of the scattering system. The gray shaded regions correspond to left and right leads connected to the scattering region shaded in blue. (b) Conductance for the uncoupled (blue curve) and coupled (black lines) systems. Vertical dashed gray lines mark the localized states of the molecule (the ring and one precursor). Local transmission for (c) Conductive, (d) Constricted, and (e) Resonant cases, calculated at energies marked by red dots in the corresponding insets. Insets: conductance for the uncoupled and coupled systems and localized states of the molecule in dashed vertical lines.
Setting the scattering problem as shown in Fig. 6(a), the electronic properties are obtained by considering the left and right leads composed of $N_{1}$-OQCs, and the center region formed by a piece of $N_{1}$-OQC attached to the molecule. The localized states of the molecule are represented by gray dashed vertical lines exhibited together with the conductance results for the uncoupled N1-OQC and the coupled system composed of chain+precursor+corral, represented by blue and black curves, respectively, in Fig. 6(b). The attachment of the molecule to the OQC affects the ballistic transport of the original $N_{1}$-OQC (blue curves) usually reducing the typical 1D channel of the $2e^{2}/h$ quantum flux. However, at particular energy values, more dramatic changes are achieved, such as full conduction suppression or resonant state.
When spatially coupled, molecular states and chain continuum states may coexist, giving rise to different responses, here distinguished by conductive, constricted and resonant cases, as shown in the normalized local transmission results in Fig. 6(c-e) with insets highlighting each particular physical phenomena. The local transmission results of panels (c) and (e) are examples in which the $G(E)$ values match perfectly with energy states of the molecule (red dots). However, in the resonant case, the electrons are trapped at very narrowed energy ranges, exhibiting almost ballistic transport. For the conductive case, the molecular states seem to be absorbed by the continuum of the OQC. Moreover, such local transmission exhibits a preferential central route that crosses the middle of the chain, although all bonds of the chain are effectively visited, differently from those of the corral sites that do not contribute to transmission. Interestingly, the resonant case is characterized by a local transmission that occurs in both chain and molecule, due to a small but sufficient percolation wave probability inside the precursor connecting chain and corral (gray shades). In fact, the sign is amplified at the internal part of the coupled corral indicated by purple in the color bar, as shown in Fig. 6(e).
Generally speaking, several localized states of the molecule have trapped waves into the continuous band of the coupled system (molecule+chain). In the coupled regime, emergent resonant states of the finite system ($E_{res}$) may lead to highly localized coupled modes within the continuum states, resulting in bound states in the continuum (BIC). To measure the intensity of this sign in electronic devices, we define a dimensionless quantity, $\mathcal{Q}=\frac{|E_{res}|}{\Delta E_{res}}$, where the strength of this interaction is related to the Fano antiresonance width $\Delta E_{res}$. The $\mathcal{Q}$ factor provides a measure of the quality of the resonant state. It directly quantifies spectral broadening and hybridization, indicating the extent to which the sharp, discrete energy levels of the isolated molecule are absorbed into the continuous spectrum of the chain upon coupling. By definition, true BICs are related to $\Delta E_{res}\rightarrow 0$, and consequently $\mathcal{Q}\rightarrow\infty$ [47]. Considering the quality factor for the present resonant example [Fig. 6(e)], we notice that the coupled state, in the black curve, almost fulfills the condition $\Delta E_{res}\rightarrow 0$, being defined as a quasi-BIC state. The molecular state, in gray, is responsible for the resonance peak (black curve) corresponding to $G(E)$ close to one quantum conductance flux, similarly to the continuum states of the chain (blue lines).
For some energies of the injected current, zero conductive is provided by electronic destructive interference. As shown in the constricted case in Fig. 6(d), the state does not fully percolate through the central part, being spatially trapped between the left lead and the molecule.
4 Conclusions
In this work, we have presented a theoretical analysis of experimentally synthesized organic quantum chains (OQCs), combining density functional theory (DFT) and molecular dynamics (MD) to obtain relaxed structures. The electronic properties were accurately captured by a tight-binding (TB) model with an exponential hopping function within the ab initio relaxed structures, being in good agreement with DFT calculations.
A key consequence of the quasi-one-dimensional character of these systems is the formation of singular hierarchical conductance channels within the electronic spectra. The confinement is enhanced by inserting multiple precursors within the unit cell, leading to a partitioning of continuous states, creating a hierarchy of pseudo-gaps. This electronic structure makes these materials highly promising for multi-band light-emitting devices, as it promotes controlled electronic transitions between the emergent sub-bands. Furthermore, the molecular chains, are characterized by an scale invariant large energy gap of $\Delta E=2.0$ eV for all proposed $N_{i}$ OQCs. This prediction matches very well with the measured experimental value [1].
Our analysis also revealed distinct localization phenomena. Compact localized states (CLS) were identified at fundamental energy levels, where the probability density is confined to a small number of sites within the unit cell. Additionally, other robust localized states, manifested as very narrow conductance channels, were observed to persist regardless the increasing number of precursors. Finally, we propose a quantum device based on the experimental formations of the carbon structures. The system composed of a OQC attached to a quantum corral shows prominent interferences and bound states in the continuum in the conductance results. For conductive situations, the transmission probability occurs in the very center of the OQC, being completely suppressed along the coupled corral. Resonant cases appear as quasi-BICs with transmission occurring along both chain and molecule, being amplified inside the corral.
We conclude that such systems are promising building blocks for the construction of carbon-based nano devices, exhibiting a combination of robust structural stability with highly tunable electronic properties. In particular, it has been possible to engineer a hierarchy of pseudo-gaps and induce localized states through smart molecular devices design.
Acknowledgements
The authors thank the INCT de Nanomateriais de Carbono for providing support on the computational infrastructure. LLL thanks the National Council for Scientific and Technological Development (CNPq) scholarship. AL thanks FAPERJ under grant E-26/200.569/2023. D.S.G. acknowledges financial support from the Coordination for the Improvement of Higher Education Personnel (CAPES) under grant No. 88887.102348/2025-00 (Finance Code 001). M.L.P.J. acknowledges financial support from the Federal District Research Support Foundation (FAPDF) under grant No. 00193-00001807/2023-16 and from the CNPq under grant No. 444921/2024-9. The authors thank P. Orellana for fruitful discussions.
Appendix A Structure and Energy of OQCs
With respect to bond lengths obtained by the methods described in Section II, all $\text{N}_{i}$-OQC systems display similar structural features regardless of the chain length. The systems consist exclusively of carbon and hydrogen atoms forming sp2 hybridizations with five- and six-fold carbon rings, the edges of which are passivated by hydrogen atoms. In pentagonal rings, the bond lengths C – C are approximately 1.47 Å. In hexagons, these lengths range from 1.39 Å to 1.43 Å. The inter-unit C – C bonds connecting the molecular precursors measure approximately 1.49 Å. The C – H bonds exhibit the typical value of 1.10 Å, and are not included in TB computations.
Table A1: Unit cell length $l_{z}$ for Ni-OQC systems.
[…]
N1 $19.4$
N2 $37.4$
N3 $54.8$
N4 $65.3$
N5 $72.2$
N6 $73.0$
N7 $68.8$
N8 $61.3$
N9 $61.8$
N10 $50.4$
More details on the length of the unit cell are shown in Table A1, where an increase in size is observed for i=16i=1-6, and a decrease for $i>6$ chains. This result underscores the relationship between chain-type and corral-type topologies. The argument of enhanced stability for the $N_{12}$ organic quantum corral proposed by Peng et al. [1] is fundamentally linked to its geometry. In the case of the open OQC structures considered here, the threshold at $i=6$ aligns with the corral configuration, as the systems begin to expand laterally and display an increased curvature beyond this point. Nearly identical $l_{z}$ values are observed for $i=8,9$, followed by a significant reduction in $l_{z}$ for $i=10$, when inter-chain interactions become significant.
From an energetic perspective, the cohesive energy of the $\text{N}_{i}$-OQC systems was calculated according to the expression,
$E_{\text{coh}}(\text{N}_{i}\text{-OQC})=\frac{E_{\text{T}}(\text{N}_{i}\text{-OQC})-40iE_{\text{C}}-20iE_{\text{H}}}{60i}$ (8)
where $E_{\text{T}}(\text{N}_{i}\text{-OQC})$ is the total energy of the $\text{N}_{i}$-OQC configuration, and $E_{\text{C}}$ and $E_{\text{H}}$ are the total energies of the isolated carbon and hydrogen atoms, respectively. This formulation normalizes the total energy per atom, enabling a direct comparison across different system sizes and topologies.
For all configurations investigated in this work, in contrast to the circular topologies reported in Ref. [1], the cohesive energy remains constant, with $E_{\text{coh}}(\text{N}_{i}\text{-OQC})=\text{-7.99}$ eV/atom, regardless of the value of $i$. In the case of quantum corrals, the minimal cohesive energy was reported for $N_{12}$, which corresponds to $N_{6}$-OQC in the present notation. This difference arises from the topological distinctions between open-chain and closed-loop geometries. The open-chain structures studied here are not subjected to the same geometric constraints that force the molecules into less favorable conformations because the chains are free to adapt their curvature while maintaining energetically equivalent configurations. Furthermore, our investigation extends up to $i=10$, which corresponds to the limit where chain-chain begin to occur, via H-H interactions, within the OQC structure, as illustrated in Fig. 1(m). For values of $i$ greater than ten, additional topological constraints are expected to impact the stability and morphology of the system.